# AIR-HLoc: Adaptive Retrieved Images Selection for Efficient Visual Localisation

### v2

@@ Abstract
State-of-the-art hierarchical localisation pipelines (HLoc) employ image retrieval (IR) to establish 2D-3D correspondences by selecting the top-$k$ most similar images from a reference database. While increasing $k$ improves localisation robustness, it also linearly increases computational cost and runtime, creating a significant bottleneck. This paper investigates the relationship between global and local descriptors, showing that greater similarity between the global descriptors of query and database images increases the proportion of feature matches. Low similarity queries significantly benefit from increasing $k$, while high similarity queries rapidly experience diminishing returns. Building on these observations, we propose an adaptive strategy that adjusts $k$ based on the similarity between the query’s global descriptor and those in the database, effectively mitigating the feature-matching bottleneck. Our approach optimizes processing time without sacrificing accuracy. Experiments on three indoor and outdoor datasets show that AIR-HLoc reduces feature matching time by up to 30%, while preserving state-of-the-art accuracy. The results demonstrate that AIR-HLoc facilitates a latency-sensitive localisation system.
@@ I INTRODUCTION
Visual localisation systems estimate the 6 degrees of freedom (6DoF) absolute camera pose for query images. Accurate and real-time absolute 6DoF pose estimation is crucial in diverse applications, from augmented reality to navigation and decision-making for mobile robots. Standard structure-based methods [1, 2, 3, 4, 5, 6, 7] achieve high accuracy by detecting and matching visual features in query images with a 3D representation of the environment. To reduce the search space in large datasets, standard hierarchical localisation pipelines (HLoc) [3, 7] include an image retrieval (IR) module to establish 2D–2D matches between the query image and the top-$k$ ranked database images. The $k$ retrieved images are used to produce 2D-3D correspondences based on a smaller subset of the environment 3D model. Camera pose is finally estimated via RANSAC and Perspective-n-Point (PnP) algorithms [8, 9].
@@
Figure 1: AIR-HLoc proposes a practical but efficient solution to increase localisation efficiency while maintaining accuracy as a structure-based visual localisation framework.
@@
These methods typically use a fixed value of $k$, determined empirically to balance robustness and performance for a given dataset [10]. The choice of $k$ is crucial, as it directly impacts localisation accuracy and runtime. This work seeks to optimize $k$ selection to minimize runtime without compromising localisation accuracy. A larger $k$ increases the number of potential matches from similar images, enhancing localisation robustness for challenging queries and improving overall positioning accuracy. However, it also extends feature matching time, becoming a key bottleneck in the pipeline, and may introduce noisy matches. Conversely, a smaller $k$ reduces runtime and improves efficiency, but at the cost of localisation accuracy. Previous works [10, 6] recommend using a moderate value of $k$, typically between 5 and 30. However, applying a fixed $k$ for all queries can reduce accuracy (if $k$ is too small) or waste computational resources and processing time (if $k$ is too large).
@@
To address this challenge, we explore strategies for achieving a better trade-off by selecting an appropriate $k$ for each query in real-world latency-sensitive visual localisation systems, as shown in Figure 1. Since global descriptors used in IR methods provide a high-level image representation, we calculate the cosine similarity between the global descriptors of the query image and the top-$3$ retrievals. This serves as the basis for the scoring mechanism described in Section III-C, which quantifies the query image’s similarity to the reference database images. Our observations indicate that query images with varying scores exhibit different levels of localisation difficulty. Higher scores, representing greater similarity and more local feature matches, correspond to easier localisation, where fewer images are required to establish sufficient 2D-3D correspondences.
Building on this insight, we introduce an adaptive retrieved image selection (AIR) approach for the hierarchical localisation pipeline, AIR-HLoc, which adaptively adjusts the number of retrieved images based on these scores for different queries. This method reduces the number of retrieved images and feature matching cost while maintaining localisation accuracy.
@@
    We propose using the cosine similarity of global descriptors to assess the localisation difficulty of queries. Building on this, we introduce an adaptive retrieved images selection (AIR) approach for HLoc. For easier queries, we retrieve fewer images, while for more challenging queries, we retrieve more images. Additionally, we present a novel metric, retrieve accuracy yield ratio (RAYR), $\zeta$, to quantify the improvement in localisation accuracy as $k$ varies across different queries. This insight facilitates more informed $k$ selection and algorithm design in practical applications.
@@
    Our proposed system, AIR-HLoc, reduces the average matching time by 30%, 26%, and 11% across Cambridge Landmarks, 7Scenes, and Aachen Day-Night-v1.1 datasets, respectively, while maintaining state-of-the-art accuracy. On the Cambridge dataset, AIR-HLoc reduces the average query runtime by 2433 ms at $k=10$ on the Jetson Orin platform, while achieving higher accuracy than HLoc.
@@ II-B Image Retrieval
Image Retrieval (IR) is commonly used for place recognition and visual localisation tasks. IR fetches top $k$ relevant and similar database images given query images by extracting global descriptor from pre-trained models [22, 24, 35, 36]. We focus on IR in structure-based localisation in this paper.  [10] studies the performance correlation between classical place recognition and localisation tasks and finds almost no improvement in visual localisation when $k$ is larger than 20. Mobile sensors’ data are utilized to reduce the database’s research space when doing IR in [37][38, 6] use a fusion of global descriptors in the visual localisation task. This study aims to improve the speed of HLoc by adaptively allocating the number of retrieved images for more efficient matching and pose estimation instead of using a constant $k$ for all queries in previous work.
@@ III Methods
Given a query image $I^{q}\in\mathbb{R}^{H\times W\times 3}$, HLoc predicts a 6DoF camera pose $\hat{p}_{k}=[\mathbf{\hat{x}}_{k},\mathbf{\hat{q}}_{k}]$ by retrieving the top-$k$ similar images to $I^{q}$. $\mathbf{\hat{x}}_{k}\in\mathbb{R}^{3}$ denotes the estimated global translation, $\mathbf{\hat{q}}_{k}\in\mathbb{R}^{4}$ denotes the estimated quaternion, which encodes the rotation. In this section, we analyze the impact of the hyperparameter $k$ on the runtime and localisation accuracy for different query images $I^{q}$. Based on our findings, we propose a novel scheme, AIR-HLoc, to enhance the efficiency of HLoc while maintaining the accuracy.
@@ III-A Problem Formulation
Feature matching is the bottleneck of HLoc. The larger the $k$, the longer it takes to feature matching. We show the detailed data in Section IV-F. We aim to reduce the feature-matching cost by retrieving top-$k^{*}$ reference images, $k^{*}\leq k$ for different query image $I^{q}$ while maintaining robustness and accuracy. We first investigate several mainstream IR methods, AP-GeM [24], NetVLAD [22], and EigenPlaces [35], to show the local descriptors matching is highly corresponding to global visual similarity between query and reference database images and different queries have different benefits for accuracy when increasing the number of $k$. So we argue that existing methods retrieve an equal number of images for all queries is a waste of processing time and computation power.
@@ III-B Relationship between match ratio and similarity
Given a query image $I^{q}$, the image database is $D$. An IR model $E$ extracts the global descriptors of the query image $I^{q}$ as $g_{q}$. $E$ extracts the global descriptors of reference images $I^{r},I^{r}\in D$ as $g_{r}$. The similarity between $I^{q}$ and $I^{r}$ is
$\cos(g^{q},g^{r})=\frac{g^{q}\cdot g^{r}}{||g^{q}||_{2}\cdot||g^{r}||_{2}}$ (1)
Similarly, a local feature extraction model $F$ extracts the local feature of the query image $I^{q}$ as $f_{q}$. $F$ extracts the local feature of $I^{r}$ as $f^{r}$. The number of local feature points of $I^{q}$ is $N(f^{q})$, and the number of matched local features between $f^{r}$ and $f^{q}$ is $M(f^{q},f^{r})$. We define the match ratio as:
match $\text{{match ratio}}=\frac{M(f^{q},f^{r})}{N(f^{q})}$ (2)
@@
Figure 2 shows the strong correlation between cosine similarity in Equation 1 and match ratio, as described in Equation 2, for each $I^{q}$ and $I^{r}$ pair. The number in parentheses of each subplot is the Pearson correlation coefficient (PCC) [39] and Spearman correlation coefficient (SRC) [40] ranges from -1 to 1. A larger PCC value indicates a stronger linear positive correlation, while SRC values close to 1 suggest a strong positive monotonic relationship. We opt to use the match ratio instead of the absolute number of matches in our analysis, as some query images exhibit textureless regions with fewer feature points, yielding only a few matches even when highly similar images are retrieved. Utilizing a match ratio from 0 to 1 provides a more meaningful measure. Our analysis incorporates three distinct IR models (AP-GeM, NetVLAD, and EigenPlaces). AP-GeM and NetVLAD serve as representative IR models in the visual localisation pipeline [6, 3, 10]. A strong positive correlation between match ratio and similarity is observed for all three IR models across three different datasets. Therefore, if a query image can find very similar images, then a small number of retrieved images can contribute enough matches for pose estimation. For queries that have difficulty finding similar images, we need more retrieved images.
@@ III-C Adaptive Retrieved Images Selection
According to the correlation observed in Figure 2, we propose to use the similarity between the images in the database and the query to evaluate the localisation difficulty of query images. We first store the global descriptors $G^{r}$ of all the images in the reference images database offline. For a query image $I^{q}$, we get the global descriptor $g^{q}$ using a pre-train IR model. Then, we calculate the cosine similarity between $g^{q}$ and each $g^{r}\in G^{r}$. The similarity is then ranked from highest to lowest, with the mean of the top $n$ being the score $S(I^{q})$ of the $I^{q}$.
$S(I^{q})=\frac{1}{n}\sum\limits_{j\in J}\cos(g^{q},g^{j})$ (3)
@@
III-D Retrieve Accuracy Yield Ratio
To assess the impact of increasing $k$ on localisation accuracy, we propose a new metric called the Retrieve Accuracy Yield Ratio (RAYR), denoted as $\zeta=(\zeta_{T}(k),\zeta_{R}(k))$. The term $\zeta_{T}(k)$ represents the reduction in translation error relative to the baseline $k=1$ for a given query image $I^{q}$, while $\zeta_{R}(k)$ captures the reduction in rotation error. Higher values of $\zeta_{T}(k)$ or $\zeta_{R}(k)$ signify a greater average contribution of the retrieved top-$k$ images to the overall improvement in localisation accuracy.
$\zeta_{T}(k)=-\frac{\text{ATE}(k)-\text{ATE}(1)}{k-1},\quad k\geq 2$ (4)
$\zeta_{R}(k)=-\frac{\text{ARE}(k)-\text{ARE}(1)}{k-1},\quad k\geq 2$ (5)
@@ IV-A Datasets
To fully demonstrate the effectiveness of our approach, we selected three popular visual localisation datasets that encompass indoor scenes, large-scale outdoor scenes, and challenges involving day-night changes, moving objects, and motion blur. The 7Scenes dataset [43, 44] is an indoor dataset consisting of seven small scenes from $1m^{3}$ to $18m^{3}$. We use the rendered depth maps provided by [15] to improve the accuracy of the sparse point cloud. The Cambridge Landmarks [25] is a large-scale outdoor dataset with five scenes. The Aachen Day-Night-v1.1 [41, 42] comprises 6697 training and 1015 test images taken from the Aachen, Germany. The database images were captured during the daytime using handheld cameras, while query images were taken with three mobile phones during both day and night. The test set is divided into two conditions: day and night. Nighttime images are only for testing.
@@ IV-B Implementation details
We integrate NetVLAD and EigenPlaces into AIR-HLoc and conduct visual localisation experiments because these two IR models have stronger correlation with the match ratio than AP-GeM as shown in Figure 2. Different IR methods have an impact on similarity thresholds. As shown in Figure 2, although the correlation between the three IR methods is similar for each dataset, the range distribution of similarity is slightly different. For Eigenplaces, the distribution of similarity on all datasets is 0.3 to 0.95, but NetVLAD and AP-GeM are not so uniform. For NetVLAD in particular, the range of similarity on Aachen Day-Night-v1.1 dataset is smaller than on Cambridge Landmarks and 7Scenes datasets. There should be two reasons for this. Firstly, the Aachen-V1.1 dataset represents a handheld motion pattern for AR applications in city-scale environments, and Eigenplaces is more robust to camera viewpoint changes than NetVLAD and AP-GeM [35]. Secondly, the Aachen Day-Night-v1.1 dataset has the cross-device domain gap that reference images and query images have different resolutions and camera intrinsic parameters, while images of 7scenes and Cambridge datasets were captured with the same camera.
Therefore, for NetVLAD, we set $\gamma^{l}=0.4$ and $\gamma^{h}=0.6$ for Cambridge and 7Scenes datasets, $\gamma^{l}=0.3$ and $\gamma^{h}=0.5$ for Aachen-V1.1 dataset. For EigenPlaces, We set $\gamma^{l}=0.7$ and $\gamma^{h}=0.9$ for all three datasets. We set $\alpha=0.5$ and $\beta=0.7$ for all IR methods and all three datasets. We conduct experiment based on the HLoc toolbox [3].

### v3

@@ Abstract
State-of-the-art hierarchical localisation pipelines (HLoc) employ image retrieval (IR) to establish 2D-3D correspondences by selecting the top-$k$ most similar images from a reference database. While increasing $k$ improves localisation robustness, it also linearly increases computational cost and runtime, creating a significant bottleneck. This paper investigates the relationship between global and local descriptors, showing that greater similarity between the global descriptors of query and database images increases the proportion of feature matches. Low similarity queries significantly benefit from increasing $k$, while high similarity queries rapidly experience diminishing returns. Building on these observations, we propose an adaptive strategy that adjusts $k$ based on the similarity between the query’s global descriptor and those in the database, effectively mitigating the feature-matching bottleneck. Our approach reduces computational costs and processing time without sacrificing accuracy. Experiments on three indoor and outdoor datasets show that AIR-HLoc reduces feature matching time by up to 30% while preserving state-of-the-art accuracy. The results demonstrate that AIR-HLoc facilitates a latency-sensitive localisation system.
@@ I INTRODUCTION
Visual localisation systems estimate 6 degrees of freedom (6DoF) absolute camera poses for query images. Accurate and real-time absolute 6DoF pose estimation is crucial in diverse applications, from augmented reality to navigation and decision-making for mobile robots. Standard structure-based methods [1, 2, 3, 4, 5, 6, 7] achieve high accuracy by detecting and matching visual features in query images with a 3D representation of the environment. To reduce the search space in large datasets, standard hierarchical localisation pipelines (HLoc) [3, 7] include an image retrieval (IR) module to establish 2D–2D matches between the query image and the top-$k$ ranked database images. The $k$ retrieved images are used to build 2D-3D correspondences based on a smaller subset of the environment’s 3D model. Camera pose is finally estimated via RANSAC and Perspective-n-Point (PnP) algorithms [8, 9].
@@
Figure 1: AIR-HLoc proposes an efficient yet practical solution to enhance localisation efficiency while maintaining accuracy by adaptively retrieving a varying number of images for different queries.
@@
These methods typically use a fixed value of $k$, determined empirically to balance robustness and performance for a given dataset [10]. The choice of $k$ is crucial, as it directly impacts localisation accuracy and runtime. This work seeks to optimize $k$ selection to minimize runtime without compromising localisation accuracy A larger $k$ increases the number of potential matches from similar images, enhancing localisation robustness for challenging queries and improving overall positioning accuracy. However, it also extends feature matching time, becoming a key bottleneck in the pipeline, and may introduce noisy matches. Conversely, a smaller $k$ reduces runtime and improves efficiency, but at the cost of localisation accuracy. Previous works [10, 6] recommend using a moderate value of $k$, typically between 5 and 30. However, applying a fixed $k$ for all queries can reduce accuracy (if $k$ is too small) or waste computational resources and processing time (if $k$ is too large).
@@
To address this challenge, we explore strategies for achieving a better trade-off by selecting an appropriate $k$ for each query in real-world latency-sensitive visual localisation systems, as shown in Figure 1. Since global descriptors used in IR methods provide a high-level image representation, we calculate the cosine similarity between the global descriptors of the query image and the top-$3$ retrievals. This serves as the basis for the scoring mechanism described in Section III-C, which quantifies the query image’s similarity to the reference database images. Our observations indicate that query images with varying scores exhibit different levels of localisation difficulty. Higher scores, representing greater similarity and more local feature matches, correspond to easier localisation, where fewer images are required to establish sufficient 2D-3D correspondences. we introduce an adaptive retrieved image selectionp two benefits. Firsthe, it allows us to resrieve fewer images while, which adaptivelocadjust accurnd, thore of each query reflects the uncertainty for different queries of posmethod reduces the number ofr of retrdimages and feature matching costsfohile maintain of differsnt difficultie.
@@
    We propose using the cosine similarity of global descriptors to assess the localisation difficulty of queries. Building on this, we introduce an adaptive retrieved image selection (AIR) approach for HLoc. For easier queries, we retrieve fewer images, while for more challenging queries, we retrieve more images. Additionally, we present a novel metric, Mean Localisation Improvement Per-retrieved-image (MLIP), $\zeta$, to quantify the improvement in localisation accuracy as $k$ varies across different queries. This insight facilitates more informed $k$ selection and algorithm design in practical applications.
@@
    Our AIR-HLoc reduces the average matching time by 30%, 26%, and 11% across Cambridge Landmarks, 7Scenes, and Aachen Day-Night-v1.1 datasets, respectively, while maintaining state-of-the-art accuracy. On the Cambridge dataset, AIR-HLoc reduces the average query runtime by 2433 ms at $k=10$ on the Jetson Orin platform, while achieving higher accuracy than HLoc.
@@ II-B Image Retrieval
Image Retrieval (IR) is commonly used for place recognition and visual localisation tasks. IR fetches top-$k$ relevant and similar database images given query images by extracting global descriptors from pre-trained models [22, 24, 35, 36]. We focus on IR in structure-based localisation in this paper.  [10] studies the performance correlation between classical place recognition and localisation tasks and finds almost no improvement in visual localisation when $k$ is larger than 20. Mobile sensors’ data are utilized to reduce the database’s research space when doing IR in [37][38, 6] use a fusion of global descriptors in the visual localisation task. This study aims to reduce computational costs and processing time of HLoc by adaptively allocating the number of retrieved images for more efficient matching instead of using a constant $k$ for all queries in previous work.
@@ III Methods
Given a query image $I^{q}\in\mathbb{R}^{H\times W\times 3}$, HLoc predicts a 6DoF camera pose $\hat{p}_{k}=[\mathbf{\hat{x}}_{k},\mathbf{\hat{q}}_{k}]$ by retrieving the top-$k$ similar images to $I^{q}$. $\mathbf{\hat{x}}_{k}\in\mathbb{R}^{3}$ denotes the estimated global translation, $\mathbf{\hat{q}}_{k}\in\mathbb{R}^{4}$ denotes the estimated quaternion, which encodes the rotation. In this section, we analyze the impact of the hyperparameter $k$ on the runtime and localisation accuracy for different query images. Based on our findings, we propose a simple yet effective scheme, AIR-HLoc, to enhance the efficiency of HLoc while maintaining the accuracy.
@@ III-A Problem Formulation
Feature matching is the bottleneck of Hloc. The larger the $k$, the longer it takes to perform feature matching. We show the detailed data in Section IV-F. We aim to reduce the feature-matching cost by retrieving top-$k^{*}$ reference images, $k^{*}\leq k$ for the different query image $I^{q}$ while maintaining robustness and accuracy. We first investigate several mainstream image retrieval methods, including AP-GeM [24], NetVLAD [22], and EigenPlaces [35]. We aim to show that feature matching strongly correlates with the global visual similarity between query and reference database images. Furthermore, different queries benefit to varying degrees from increasing $k$ in terms of accuracy. So we argue that existing methods of retrieving an equal number of images for all queries are a waste of time and computational power.
@@ III-B Relationship between match ratio and similarity
Given a query image $I^{q}$, the image database is $D$. An IR model $E$ extracts the global descriptors of the query image $I^{q}$ as $g^{q}$. $E$ extracts the global descriptors of reference images $I^{r},I^{r}\in D$ as $g^{r}$. The similarity between $I^{q}$ and $I^{r}$ is
$\cos(g^{q},g^{r})=\frac{g^{q}\cdot g^{r}}{||g^{q}||_{2}\cdot||g^{r}||_{2}}.$ (1)
Similarly, a local feature extraction model $F$ extracts the local feature of the query image $I^{q}$ as $f^{q}$. $F$ extracts the local feature of $I^{r}$ as $f^{r}$. The number of local feature points of $I^{q}$ is $N(f^{q})$, and the number of matched local features between $f^{r}$ and $f^{q}$ is $M(f^{q},f^{r})$. We define the match ratio as:
match $\text{{match ratio}}=\frac{M(f^{q},f^{r})}{N(f^{q})},$ (2)
@@
Figure 2 shows the strong correlation between cosine similarity in Equation 1 and match ratio, as described in Equation 2, for each $I^{q}$ and $I^{r}$ pair. The number in parentheses of each subplot is the Pearson correlation coefficient (PCC) [39] and Spearman correlation coefficient (SRC) [40], both ranging from -1 to 1. A larger PCC value indicates a stronger linear positive correlation, while an SRC value close to 1 suggests a strong positive monotonic relationship. We opt to use the match ratio instead of the absolute number of matches in our analysis, as some query images exhibit textureless regions with fewer feature points, yielding only a few matches even when highly similar images are retrieved. Utilizing a match ratio from 0 to 1 provides a more meaningful measure. Our analysis incorporates three distinct IR models (AP-GeM, NetVLAD, and EigenPlaces). AP-GeM and NetVLAD serve as representative IR models in the visual localisation pipeline [6, 3, 10]. A strong positive correlation between match ratio and similarity is observed for all three IR models across three different datasets. Therefore, if a query image can find very similar images, then a small number of retrieved images can contribute enough matches for pose estimation. For queries that have difficulty finding similar images, we need more retrieved images.
@@ III-C Adaptive Retrieved Images Selection
According to the correlation observed in Figure 2, we propose to use the similarity between the images in the database and the query to evaluate the localisation difficulty of query images. We first store the global descriptors $G^{r}$ of all the images in the reference images database offline. For a query image $I^{q}$, we get the global descriptor $g^{q}$ using a pre-trained IR model. Then, we calculate the cosine similarity between $g^{q}$ and each $g^{r}\in G^{r}$. The similarity is then ranked from highest to lowest, with the mean of the top-$n$ being the score $S(I^{q})$ of the $I^{q}$.
$S(I^{q})=\frac{1}{n}\sum\limits_{j\in J}\cos(g^{q},g^{j}),$ (3)
@@
III-D Mean Localisation Improvement Per-retrieved-image
To assess the impact of increasing $k$ on localisation accuracy, we propose a new metric called the Mean Localisation Improvement Per-retrieved-image (MLIP), denoted as $\zeta=(\zeta_{T}(k),\zeta_{R}(k))$. The term $\zeta_{T}(k)$ represents the reduction in translation error relative to the baseline $k=1$ for a given query image $I^{q}$, while $\zeta_{R}(k)$ captures the reduction in rotation error. Higher values of $\zeta_{T}(k)$ or $\zeta_{R}(k)$ signify a greater average contribution of the retrieved top-$k$ images to the overall improvement in localisation accuracy.
$\zeta_{T}(k)=-\frac{\text{ATE}(k)-\text{ATE}(1)}{k-1},\quad k\geq 2,$ (4)
$\zeta_{R}(k)=-\frac{\text{ARE}(k)-\text{ARE}(1)}{k-1},\quad k\geq 2.$ (5)
@@ IV-A Datasets
To fully demonstrate the effectiveness of our approach, we selected three popular visual localisation datasets that encompass indoor scenes, large-scale outdoor scenes, and challenges involving day-night changes, moving objects, and motion blur. The 7Scenes dataset [43, 44] is an indoor dataset consisting of seven small scenes from $1m^{3}$ to $18m^{3}$. The Cambridge Landmarks [25] dataset is a large-scale outdoor dataset with five scenes. The Aachen Day-Night-v1.1 [41, 42] dataset comprises 6697 training and 1015 test images taken from Aachen, Germany. The database images were captured during the daytime using handheld cameras, while query images were collected with three different mobile phones in both day and night conditions. The test set is split into two subsets: day and night. Nighttime images are only for testing.
@@ IV-B Implementation details
We integrate NetVLAD (NV) and EigenPlaces (EP) into AIR-HLoc and conduct visual localisation experiments because these two IR models have a stronger correlation with the match ratio than AP-GeM, as shown in Figure 2. Different IR methods have an impact on similarity thresholds. As shown in Figure 2, although the correlation between the three IR methods is similar for each dataset, the range distribution of similarity is slightly different. For EigenPlaces, the distribution of similarity on all datasets is 0.3 to 0.95, but NetVLAD and AP-GeM are not so uniform. For NetVLAD in particular, the range of similarity on the Aachen Day-Night-v1.1 dataset is smaller than on the Cambridge Landmarks and 7Scenes datasets. There are two possible reasons for this phenomenon. Firstly, the Aachen-V1.1 dataset represents a handheld motion pattern for AR applications in large city-scale environments, and EigenPlaces is more robust to camera viewpoint changes than NetVLAD and AP-GeM [35]. Secondly, the Aachen Day-Night-v1.1 dataset has a cross-device domain gap, as the reference and query images have different resolutions and camera intrinsic parameters, whereas the images in the 7Scenes and Cambridge datasets were captured with the same camera.
Therefore, for NetVLAD, we set $\gamma^{l}=0.4$ and $\gamma^{h}=0.6$ for Cambridge and 7Scenes datasets, $\gamma^{l}=0.3$ and $\gamma^{h}=0.5$ for the Aachen-V1.1 dataset. For EigenPlaces, we set $\gamma^{l}=0.7$ and $\gamma^{h}=0.9$ for all three datasets. We set $\alpha=0.5$ and $\beta=0.7$ for all IR methods and all three datasets. We conduct experiments based on the HLoc toolbox [3].